OpenStudy (anonymous):

4 friends have different numbers of cows on their farms. Jorge has 3 times as many cows as carla. Marcus has half as many cows as jorge. Dominque has 2 less than 4 times as many cows as carla. how many cows does each person have, total is 169?

OpenStudy (anonymous):

Let Carla as x So, Jorge = 3x Marcus = 3x/2 Dominique = 4c - 2 Total = 169 Carla + Jorge + Marcus + Dominique = total x + 3x + 3x/2 + 4x - 2 = 169 8x + 3x/2 = 171 16x + 3x = 342 19x = 342 x = 342/19 x = 18 Carla = x = 18 Jorge = 3x = 54 Marcus = 3x/2 = 54/2 = 27 Dominique = 4x - 2 = 72 - 2 = 70 Check (finding the total number of cows which is 169 18 + 54 + 27 + 70 = 169

OpenStudy (anonymous):

TY

OpenStudy (anonymous):

np